OpenStudy (anonymous):

Solve the differential equation y' = x^2 + x^(-1), given y(-1) = 2.

terenzreignz (terenzreignz):

Pretty much this, then: \[\Large \frac{dy}{dx} = x^2 + \frac1x\] ?

terenzreignz (terenzreignz):

@creampuffu wake up lol

OpenStudy (anonymous):

Yeah, that's what it looks like?

OpenStudy (anonymous):

Do I evaluate for the antiderivative then solve for C?

terenzreignz (terenzreignz):

Pretty much, yes :D ^_^\[\Large \int \frac{dy}{dx}dx = \int \left(x^2 + \frac1x\right)dx\]\[\Large \int \frac{dy}{\cancel{dx}}\cancel{dx} = \int \left(x^2 + \frac1x\right)dx\]\[\Large \int dy = \int \left(x^2 + \frac1x\right)dx\]

OpenStudy (anonymous):

\[\frac{ x^3 }{ 3 }+lnx + c\] then -1/3 +ln(1) + c = 2? c=7/3

terenzreignz (terenzreignz):

Yeah, that's right. You sure that you actually needed help here? XD By the way, just to avoid those nasty nitpicking instructors, \[\Large \int \frac1x dx = \ln|x| + C\] don't forget those absolute value bars, actually, or else ln(-1) won't make sense ^_^ But other than that, good job :)